# Beyond the Visible: Multispectral Vision-Language Learning for Earth Observation

### v1

@@ 6 Results
Table 1: Evaluation results on EuroSAT, BigEarthNet, METER-ML, and the overall average. We report zero-shot classification results in accuracy (%) $\uparrow$ and text-to-image retrieval results in mAP@100 (%) $\uparrow$. The best-performing model is highlighted in bold and the second-best model is underlined.
@@
We provide example predictions from Llama3-MS-CLIP in Figure 5 to illustrate common behavior. On EuroSAT, the model is nearly always correct for general classes like residential, sea/lake, or industrial, but it confuses certain pairs such as permanent crop and annual crop. On METER-ML, samples are often mistakenly classified as wastewater treatment plants, while concentrated animal feedings operations (COFAs) like farms and other features (negative class) are mostly correctly identified.
@@
In BigEarthNet’s multi-label scenario, our approach tends to produce numerous false positives. However, these misclassifications are usually semantically correlated (e.g., inland waters or beaches predictions for marine waters). Introducing an extra negative class (other features) mitigates false positives but leads to much more false negatives. This results in higher accuracy overall, but the F1 score drops for six of the seven models. Since F1 reflects both precision and recall, we report the original method here and the alternative strategy in the supplementary material. In both cases, MS-CLIP achieves the highest F1 and significantly outperforms five other models in accuracy.
@@ 6.2 Ablation Studies
We conduct several ablation experiments to systematically study various design choices. Specifically, we investigate prompt templates, weight initialization for the patch embedding, the use of different input bands, and strategies for freezing/unfreezing model layers.
@@
Figure 6: Ablation experiments for the initialization of additional patch embedding channels (a) and the selected input bands (b). The final setting is in blue color.
Figure 6(a) compares two approaches to initialize the multispectral patch embeddings. While a naive solution is to set each new channel’s weights to the mean of the RGB ones, we find that zero-initialization produces superior performance in all six tasks. We hypothesize that starting from zero, together with a short warm-up phase, lets the model adjust gradually to the additional input channels instead of the more sudden interruption with mean initialization.
In Figure 6(b), we examine the effect of using three, ten, or all twelve Sentinel-2 bands. Although using all bands might seem beneficial, certain tasks decline in performance compared to using only the RGB bands. Omitting bands 1 and 10 and training on ten bands leads to the overall best results. The dropped bands both have a 60 m spatial resolution and their information therefore might not align well with the other bands. Dropping these bands is common among EO models (e.g., in [6]) suggesting low meaningful information for many use cases.
We further compare various strategies for freezing and unfreezing model layers in Table 4. Freezing layers can, in principle, reduce the risk of catastrophic forgetting. Yet, our experiments show that fully unfreezing the image and text encoders leads to the best results when adapting to the SSL4EO-S12 image-caption data. By contrast, selectively fine-tuning only the projection layer yields lower performance than the baseline.
@@ 7 Conclusion
We introduce a multispectral vision-language dataset of low-resolution Sentinel-2 data with corresponding captions. Our automated captioning strategy scales easily, removing the need for costly human annotations. In addition, SSL4EO-S12 includes aligned Sentinel-1 data and we provide Q&A pairs that can both be leveraged for multimodal learning. This dataset enabled us to train the first CLIP-like multispectral VLM. Experiments show that our model significantly improves zero-shot classification and retrieval compared to other methods, even domain-specific adaptations trained on larger datasets. We believe that the open-sourced Llama3-MS-CLIP has considerable potential, both for immediate downstream tasks and as a vision encoder for building multispectral MLLMs.

### v2

@@ 6 Results
Table 1: Evaluation results on EuroSAT, BigEarthNet, METER-ML, and the overall average. We report zero-shot classification results in accuracy (%) $\uparrow$ and text-to-image retrieval results in mAP@100 (%) $\uparrow$. The best-performing model is highlighted in bold, and the second-best model is underlined.
@@
We provide example predictions from Llama3-MS-CLIP in Figure 5 to illustrate common behavior. On EuroSAT, the model is nearly always correct for general classes like residential, sea/lake, or industrial, but it confuses certain pairs such as permanent crop and annual crop. On METER-ML, samples are often mistakenly classified as wastewater treatment plants. In contrast, concentrated animal feedings operations (COFAs) like farms and other features (negative class) are mostly correctly identified.
@@
In BigEarthNet’s multi-label scenario, our approach tends to produce numerous false positives. However, these misclassifications are usually semantically correlated (e.g., inland waters or beaches predictions for marine waters). Introducing an extra negative class (other features) mitigates false positives but leads to many more false negatives. This results in higher accuracy overall, but the F1 score drops for six of the seven models. Since F1 reflects both precision and recall, we report the original method here and the alternative strategy in the supplementary material. In both cases, MS-CLIP achieves the highest F1 and significantly outperforms five other models in accuracy.
(a) EuroSAT
(b) METER-ML
Figure 6: UMAP plot of the embedding spaces from OpenCLIP B/16 and Llama3-MS-CLIP for EuroSAT and METER-ML using the S-2 test images. The UMAP settings are 10 nearest neighbors and min distance 0.0.
We further analyse the embedding space of the single-label datasets EuroSAT and METER-ML in Figure 6 and compare Llama3-MS-CLIP with the base model. While many EuroSAT classes overlap in the OpenCLIP UMAP visualization, the clusters are more distinct for Llama3-MS-CLIP. Differentiating between classes in METER-ML is much more difficult, which is reflected in the lower accuracy and the embedding space visualization with no large clusters for any model. However, some classes like wastewater treatment or refineries and terminals form small clusters in the Llama3-MS-CLIP plot, while being more distributed for OpenCLIP.
@@ 6.2 Ablation Studies
We conduct several ablation experiments to study various design choices systematically. Specifically, we investigate prompt templates, weight initialization for the patch embedding, different input bands, and strategies for freezing/unfreezing model layers.
@@
Figure 7: Ablation experiments for the initialization of additional patch embedding channels (a) and the selected input bands (b). The final setting is in blue.
Figure 7(a) compares two approaches to initialize the multispectral patch embeddings. While a naive solution is to set each new channel’s weights to the mean of the RGB ones, we find that zero-initialization produces superior performance in all six tasks. We hypothesize that starting from zero with a short warm-up phase lets the model adjust gradually to the additional input channels instead of the more sudden interruption with mean initialization.
We analyse the weights of the patch embedding before and after continual pre-training. The absolute values of the patch embedding for the multispectral channels are much lower than for the RGB channels, showing that the model did not fully adapt to the additional input due to the limited number of weight updates with only one million samples. At the same time, changes in the multispectral channels are $\sim$2800 times higher than the changes in the RGB weights, which are adjusted by only 0.03% compared to the OpenCLIP weights. This shows the additional information Llama3-MS-CLIP leverages from the multispectral channels and highlights the need for even larger multispectral EO datasets for longer pre-training without overfitting.
In Figure 7(b), we examine the effect of using three, ten, or all twelve Sentinel-2 bands. Although using all bands might seem beneficial, certain tasks decline in performance compared to using only the RGB bands. Omitting bands 1 and 10 and training on ten bands leads to the overall best results. The dropped bands both have a 60 m spatial resolution, and their information might not align well with the other bands. Dropping these bands is common among EO models (e.g., in [7]), suggesting low meaningful information for many use cases.
We further compare various strategies for freezing and unfreezing model layers in Table 4. The patch embedding was unfrozen in every setting to adapt to the multispectral input. Freezing layers can, in principle, reduce the risk of catastrophic forgetting. Yet, our experiments show that fully unfreezing the image and text encoders leads to the best results when adapting to multispectral data. By contrast, selectively fine-tuning only the projection layer yields lower performance than the baseline. We performed a similar ablation study for Llama3-RGB-CLIP with contrary results. Fine-tuning only the projection layer resulted in the best performance. Keeping the earlier layers frozen avoids forgetting pre-trained features with RGB input, but it cannot capture the multispectral information when including the additional multispectral channels.
@@
7 Discussion and Limitations
In this work, we demonstrate the benefit of leveraging MLLMs for accelerating image captioning in order to curate datasets for subsequent vision-language model training. While MLLMs are already able to digest and caption RGB representations of remote sensing images themselves, they cannot leverage multispectral data—leaving room for specialized models for Earth observation. We address this gap by coupling the automated caption generation on RGB imagery with multi-spectral data for the Llama3-MS-CLIP. However, we acknowledge limitations when using MLLMs to generate synthetic image captions. First, we understand that there exists a risk of propagating errors or biases (e.g., in the form of hallucinations) of the MLLM further into the self-supervised models that are trained on top of the synthetic data. It will be relevant to identify such ripple effects that result from training self-supervised models on synthetically generated data of MLLMs in order to understand how errors and biases are propagated, reduced, or reinforced by downstream models. Second, we note that the existing dataset likely benefits from increased diversity, as the word count graph in the supplementary section highlights a trend of similar topics in many of the captions. Third, we advocate for considering human-in-the-loop systems during the caption generation process in settings where errors by the MLLMs are not acceptable.
Based on the synthetically generated captions, Llama3-MS-CLIP demonstrates the benefit of using multi-spectral data during pretraining. Even though we observe significant performance improvements when leveraging multi-spectral data instead of RGB data, our experiments also show that the model is potentially not yet saturated. This is indicated by comparably low weights in the patch embedding of Llama3-MS-CLIP for the non-visible channels compared to the visible RGB spectrum. Overall, this experiment reinforces our expectation that the performance of the model will further improve with longer continuous pretraining, leveraging a larger and more diverse training corpus.
Finally, we see a possibility for future research to work on integrating and merging pixel-level training strategies with the image-level training we employ in this work. This merge could improve the model’s capability to capture detailed image nuances and help differentiate between closely related classes. Pixel-level understanding might also unlock additional progress on other tasks that we did not explore in this work, including semantic segmentation and object detection.
@@ 8 Conclusion
We introduce a multispectral vision-language dataset of low-resolution, multispectral Sentinel-2 data with corresponding captions. Our automated captioning strategy scales easily, reducing the need for costly human annotations. On top of this dataset, we build Llama3-MS-CLIP, the first CLIP-like multispectral VLM. Our experiments show that our model significantly improves zero-shot classification and retrieval compared to other methods, even domain-specific adaptations trained on larger datasets. We see significant potential in leveraging the open-sourced Llama3-MS-CLIP in downstream applications and as a vision encoder for building multispectral MLLMs.
@@
8.0.1 Acknowledgements
We thank the remote sensing experts for reviewing the generated captions and Niklas Kopp for providing the embedding space analysis.